# One-step Generative Diffusion for Realistic Extreme Image Rescaling

### v2

@@ 5.1 Comparison with JPEG Compression
5.2 Time-step Distribution
To overcome the non-uniform degradation introduced by the rescaling operation, we design a time-step alignment strategy to predict an optimal time-step $t$ based on the degraded latent variable $\hat{z}$, enabling dynamic allocation of generative capability. During the inference phase, our model can perform tiled inference, allowing it to predict different time-steps for various regions of the same image. We discuss the distribution of time-steps at both the single image and dataset scales.
Figure 17: The time-steps $t$ predicted by the TPM on a single image, along with the corresponding latent variable rescaling error map $|z-\hat{z}|$.
First, we visualize the time-steps corresponding to different regions of the image shown in Fig. 17. We also present the latent variable rescaling error map $|z-\hat{z}|$ for comparison. It can be observed that in more complex regions, the rescaling operation introduces larger errors. Correspondingly, these areas also require larger time-steps. This is because complex regions necessitate greater generative capability for restoration.
Figure 18: The distribution of time-steps predicted by the TPM on the DIV2K validation set at different rescaling scale factors. Smaller scale factor results in smaller time-steps.
Next, we present the distribution of predicted time-steps at different rescaling scales on the DIV2K validation set shown in Fig. 18. The vertical axis and the size of the bubbles represent the mean and variance of the time-steps, respectively. It can be observed that the mean of the predicted time-steps increases as the rescaling scale becomes larger. This is because larger rescaling scales result in greater image degradation, necessitating stronger generative capability for restoration. Additionally, we find that the predicted time-steps for the 16$\times$ rescaling exhibit the largest variance. This indicates that 16$\times$ rescaling is a task that balances both fidelity and generative capacity. In contrast, 8$\times$ rescaling tends to prioritize fidelity, while 32$\times$ rescaling leans more towards generative capabilities.
6 Limitations
Table 3: The performance comparison between our 4×4$\times$4 × rescaling model and VAE.
Method PSNR $\uparrow$ LPIPS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
VAE 29.00 0.1570 65.54 0.6234
Ours-4×4$\times$4 × 29.32 0.1456 65.47 0.6313
6.1 Upper Bound of OSIRDiff
Since our method performs the rescaling operation in the latent space of a pre-trained VAE, its performance ceiling is inherently constrained by the VAE itself. The VAE encoding is a lossy process and can be described as follows:
$z=\mathcal{E}(x)$ (14)
Here, $x\in R^{H\times W\times 3}$ represents original HR image and $z\in R^{\frac{H}{8}\times\frac{W}{8}\times 4}$ represents corresponding latent features. We can measure the compression rate from the perspective of the number of bits:
[…]
$=\left(H\times W\times 3\right)\times 8$ (15)
[…]
$=\left(\frac{H}{8}\times\frac{W}{8}\times 4\right)\times 16$ (16)
$\mathrm{Rate}$ $=\frac{\mathrm{Bits(z)}}{\mathrm{Bits(x)}}=\frac{1}{24}$ (17)
It can be observed that VAE encoding can be regarded as a compression process with a compression ratio of 24. From the perspective of rescaling, this is equivalent to a rescaling operation with a scale factor of $S=\sqrt{24}$. Any rescaling scale smaller than $S$ will inevitably be limited by the reconstruction performance of VAE, as the VAE encoding process has already discarded significant amounts of information.
Furthermore, we directly encode and decode input HR images using VAE without performing any rescaling operations, representing the upper performance bound of our model. Additionally, we train an OSIRDiff model for 4×4$\times$4 × rescaling and compare its performance with the reconstruction capability of the VAE, as illustrated in Table 3. It can be observed that the performance of our 4×4$\times$4 × rescaling model is very close to the reconstruction performance of the VAE, indicating that the rescaling performance at this scale is limited by the VAE reconstruction capacity. This also highlights a potential limitation of SD prior itself that it only demonstrates significant advantages when handling highly ill-posed tasks [8, 2, 46]. Consequently, our proposed OSIRDiff is similarly affected by this limitation and performs well primarily at larger rescaling scales (>$S$). In the future, we could explore the incorporation of shallow features into the VAE decoding process to enhance its fidelity, thereby enabling the OSIRDiff framework to be more applicable to small-scale rescaling tasks.
Table 4: Comparison of inference time and parameter sizes across methods.
Time (s) Total Parameters (M) Trainable Parameters (M)
IRN 0.359 34.2 34.2
HCFlow 0.363 7.24 7.24
VQIR 0.897 92.9 20.7
Ours 0.711 1352.5 63.2
6.2 Inference Time and Model Size
In Table 4, we compare the inference speed and parameter sizes of different methods. All tests are performed on an RTX 3090 using images with a resolution of $1024\times 1024$. Due to leveraging the SD architecture, our method has the largest total number of parameters and the highest number of trainable parameters. However, since our approach only requires a single denoising step, its inference time does not increase significantly compared to other methods. The inference speed of our method even surpasses that of the VQGAN-based [4] method, i.e. the VQIR. This is because, compared to VAE, the latent space of VQGAN has a higher number of channels. Additionally, the nearest-neighbor search process in VQGAN incurs significant computational overhead, adversely affecting its inference speed. In the future, we plan to explore architecture distillation techniques to further reduce the size of our image rescaling model and improve its inference speed.
@@ 7 Conclusion
We propose a novel framework called OSIRDiff to address issues such as insufficient texture and semantic inaccuracies in extreme image rescaling. Specifically, we design a DFRM to achieve bidirectional mapping between HR features and LR images. Then, we utilize a pre-trained SD model to perform a single denoising step on the rescaled features, enhancing its perceptual quality. Considering the non-uniform quality of the rescaled latent features, we propose a novel time-step alignment strategy to achieve a balance between fidelity and perceptual quality. Both qualitative and quantitative experiments demonstrate the superiority of our approach.

### v3

@@ 4.4.2 Application to Image Compression
4.4.3 Inference Efficiency Comparison
In Table 3, we provide comparisons of inference time and memory consumption for representative methods. All tests are performed on an RTX 3090 using images with a resolution of $1024\times 1024$. It can be observed that flow-based models (IRN and HCFlow) require less GPU memory consumption and enable fast inference. In contrast, StableSR, which is based on multi-step diffusion models, demands significantly more inference time and larger GPU memory consumption. Although VQIR requires less memory, the nearest neighbor search process in VQGAN results in longer inference time. For one-step diffusion-based method S3Diff, due to its degradation-guided LoRA which needs to dynamically predict LoRA parameters during inference, it requires longer ingerence time and higher memory consumption. In contrast, our proposed TADM demonstrates moderate efficiency and scalability compared to other methods. In the future, we can utilize distillation techniques to further reduce the model size.
Table 3: Comparison of inference time and memory consumption.
IRN HCFlow VQIR StableSR S3Diff Ours
Inference Time (s) 0.359 0.363 0.897 135.2 2.186 0.711
Memory Consumption (G) 0.498 0.098 0.344 6.380 5.025 2.598
4.4.4 Upper Bound of TADM
Since our method performs the rescaling operation in the latent space of a pre-trained VAE, its performance upper boud is inherently constrained by the VAE. In Table 4, we present a performance comparison between VAE and two 4$\times$ rescaling models. For VAEs, we directly feed the high-resolution images to the encoder and decoder without performing any rescaling operations. This performance represents the upper bound of our method’s performance across all rescaling factors. It can be observed that VAE significantly lags behind the latter. In the future, we could explore integrating stronger VAEs, such as those from SDXL, SD3, thereby enabling the TADM to be more applicable to small-scale rescaling tasks.
Table 4: Comparison of VAE and 4$\times$ rescaling methods.
IRN HCFlow VAE
PSNR 35.07 35.23 29.01
SSIM 0.9318 0.9346 0.8046
@@ 5 Conclusion
We propose a novel framework called TADM to address issues such as insufficient texture and semantic inaccuracies in extreme image rescaling. Specifically, we design a DFRM to achieve bidirectional mapping between HR features and LR images. Then, we utilize a pre-trained SD model to perform a single denoising step on the rescaled features, enhancing its perceptual quality. Considering the non-uniform quality of the rescaled latent features, we propose a novel time-step alignment strategy to achieve a balance between fidelity and perceptual quality. Both qualitative and quantitative experiments demonstrate the superiority of our approach.